OpenStudy (anonymous):

Form a polynomial function with a degree of 4, zeros 2-4i; 4 multiplicity 2. Please help!

OpenStudy (anonymous):

i will show u the steps then u conclude

OpenStudy (anonymous):

Perfect!!! Thank you very much!!!!!!!

OpenStudy (anonymous):

step one: you can get a zero from multiplying the complex number by its conjugant

OpenStudy (anonymous):

so in your case we have:

OpenStudy (anonymous):

\[(2-4i), (2+4i)\]

OpenStudy (anonymous):

sorry am on the phone i will continue the whole problem after am done

OpenStudy (anonymous):

No problem! I had (x-(2-4i)(x-(2+i)(x+4)^2

OpenStudy (anonymous):

yes that is correct apart from a missing 4 in ur cojugent

OpenStudy (anonymous):

so u are having a problem multiplying them

OpenStudy (anonymous):

I get to this point and I'm c lost! I don't know what to do next!

OpenStudy (anonymous):

well u have to multiply the whole thing to get the polynomial

OpenStudy (anonymous):

the best way to do that is to simplify the product let me show u

OpenStudy (anonymous):

i am writing a simple method for u to always be able to solve such a problem quickly

OpenStudy (anonymous):

\[(x+iy)(x-iy)=x^2-(iy)^2\]

OpenStudy (anonymous):

r u there?

OpenStudy (anonymous):

I'm here sorry :-)

OpenStudy (anonymous):

\[[x-(2-4i)][x-2+(2-4i)][(x+4)^2]\]

OpenStudy (anonymous):

now let's solve the complex values first

OpenStudy (anonymous):

\[=[(x-2)+4i)][(x-2)-4i)]\]

OpenStudy (anonymous):

can u see i rearranged it in terms of real and imaginary parts ok

OpenStudy (anonymous):

\[=(x-2)^2-(4i)^2\]

OpenStudy (anonymous):

Yes I follow that

OpenStudy (anonymous):

\[=x^2-4x+4+16=x^2-4x+20\]

OpenStudy (anonymous):

Oh yeah that formula you gave earlier makes sense now

OpenStudy (anonymous):

\[=[x^2-4x+20][(x+4)^2]\]

OpenStudy (anonymous):

now lets expand

OpenStudy (anonymous):

so u can see we are getting a 4th order polynomial

OpenStudy (anonymous):

\[=[x^2-4x+20][x^2+8x+16]\]

OpenStudy (anonymous):

the final answer hence is:

OpenStudy (anonymous):

\[=x^4+4x^3+4x^2+96x+320\]

OpenStudy (anonymous):

i did that on piece of paper

OpenStudy (anonymous):

I was still calculating mine, but I got the same answer!

OpenStudy (anonymous):

Thank you so very much for taking the time to actually walk me through step by step!!!!!!!

OpenStudy (anonymous):

good all the best

OpenStudy (anonymous):

don't forget my medal hehehe

OpenStudy (anonymous):

How do I do it

OpenStudy (anonymous):

how do i know hehehe, i really don't know

OpenStudy (anonymous):

i am just kidding anyways

OpenStudy (anonymous):

Figured it out :-) there you go, you have a medal !!! Thanks again!!!!!